A computer is printing out subsets of a 5 element set (possibly including the empty set).
a. at least how many sets must be printed to be sure of having at least 4 identical subsets on the list?
b. at least how many identical subsets are printed if there are 161 subsets on the list?
YOUR HELP IS GREATLY APPRECIATED!

Answer by sudhanshu_kmr(1152)


Total no. of possible subsets of 5 elements (including empty set)
= 5C0 + 5C1 + 5C2 + 5C3 +5C4 + 5C5 = 1 + 5 +10 + 10 + 5 + 1 = 32

a). no. of subset to ensure at least 4 identical subsets = 32*3 + 1

= 97



b). 161 = 32 * 5 + 1

so, at least 6 identical subsets are printed.



Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......